17 January 2008

RHIC Running:  There was a major failure in the Controls Database Server.  Had to switch over to a back-up server while the old server was rebuilt.  Beam had to be off for the switch over.

While the database access was going on, both PHENIX and STAR made accesses.  PHENIX has a bad HV mainframe that they will try to repair.

There were first attempts to inject polarized protons from the LINAC into the Booster without success.

BLIP came back on after fixing their sparking problems in Module #3. 

Physics came back on after the Maintenance Day yesterday by ~20:30.  Not too shabby!.

 


Scheduling Physicist: Mike Sivertz Last updated:
Email: sivertz@bnl.gov
Office Phone: 631-344-6102
Digital Pager: 7528 or 4300
Home Phone: 631-878-9020